Ahmer walks 795m north and 55m east 80m south find the displacement and distance?
1
Expert's answer
2020-03-09T10:55:41-0400

The distance


"d=795\:\rm m+55\:\rm m+80\:\rm m=930\:\rm m."

The displacement


"s=\sqrt{(795-80)^2+55^2}=717\:\rm m."